Sorry should have been an equal sign
e^(x^2+7x)=(1/e^12)

Answer by stanbon(75887)

Sorry should have been an equal sign
e^(x^2+7x)=(1/e^12)
------------
Cross-multiply:
e(x^2+7x+12) = 1
----
Take the natural log
x^2+7x+12 = 0
(x+3)(x+4) = 0
x = -3 or x = -4
===================
Cheers,
Stan H.